OpenStudy (anonymous):

Simplify the radical expression: √(81/121)

OpenStudy (anonymous):

 do you know what those are?

OpenStudy (anonymous):

and once you find the square roots of those values, then you have your answer!

OpenStudy (anonymous):

0.818181818181818181818181818181

OpenStudy (anonymous):

\[\sqrt{(\frac{ 81 }{ 121 })} = \sqrt{(\frac{ 9 }{ 11 })^2}=\frac{ 9 }{ 11 }\]

OpenStudy (anonymous):

is suck